# Riccardo Giordana Pozzi

### v1

@@ 6.4 Universal sectors
We now turn to the conditions that characterize the presence of universal sectors. The analysis of the explicit OPEs shows that the displacement multiplet is never exchanged, so all the requirements discussed in section 2 are satisfied. We therefore conclude that the superprimary and the displacement sectors remain decoupled up to NLO. The identification of the NLO four-point function then becomes straightforward, allowing us to bypass the standard bootstrap procedure of stating and fixing an ansatz. This case is very similar to that discussed in subsection 4.1. The four-point function of the superprimary can be obtained directly from that of the first superdescendant of the ${\mathcal{N}}=4$ CSm Wilson line superprimary [22]. This leads to
@@
In this case, the overall coefficient is fixed explicitly by the analysis in subsection 6.2 and the expansion parameter is again linked to the two point normalization as in (4.8) by
@@ 6.5 Extracting the conformal data
6.5.1 Leading order data
As discussed in section 3 the exchanged operators at leading order are those of the multiparticle sates: $\mathbb{\Lambda}\bar{\mathbb{\Lambda}}]^{L=2}_{n}$ in the chiral-antichiral channels and $\mathbb{\Lambda}\mathbb{\Lambda}]^{L=2}_{n}$ in the chiral-chiral one, all having dimensions $\Delta=3+n$. We can thus extract the leading order data by exploiting the orthogonal relation for the blocks and superblocks discussed in the appendix A.1. We thus get
@@ 6.5.1 Leading order data
6.5.2 Next-to-leading order data
We now turn to the next-to-leading order of the expansion, beginning with the chiral–chiral channel. In this case, the anomalous dimensions are extracted from the logarithmic contributions of the NLO term in the correlation function (6.15). Explicitly, they are given by
@@ 6.5.2 Next-to-leading order data
which explicitly read
@@
Exploiting the relation under the map $z\rightarrow\chi$ discussed in subsection 6.3 and the braiding relation (6.10), one can obtain the anomalous dimensions for the operators exchanged in the chiral-chiral channel. These are given by
@@
and, explicitly, the OPE coefficient read
@@ 7 The $1/3$ BPS Wilson line in ABJM
where the vertical lines stand for the action of the preserved supercharges corresponding to the considered theory and, for notation convenience, we also included the corresponding operators ($a=1,2,3$ and $r=1,2$). As in the order case, the four-point of the tilt and of the displacement (anti)chiral multiplet will be specified by the top components four-point functions of $\mathbb{R}$ ($\bar{\mathbb{R}}$) and $\mathbbm{F}$} ($\bar{\mathbbm{F}}$) respectively.
We now investigate the presence of the universal sector in both the displacement and tilt four-point functions. Focusing first on the displacement supermultiplet, one can verify exploiting the same arguments of section 3 that the displacement sector does not mix with the tilt sector and, moreover, that the individual defect operators within the multiplet do not mix among themselves. It follows that the corresponding four-point function exhibits universal features. Furthermore, since the algebraic structure coincides with that of the $1/2$ BPS line in ${\mathfrak{osp}}(4|4)$, the analysis leads to the same expressions of [22].
@@
where $C_{\mathcal{T}}$ denotes the two-point normalization of the tilt supermultiplet, which is distinct from that of the displacement multiplet. Nevertheless, if the theory is controlled by a single perturbation parameter describing deviations from the GFF, one may still expect $C_{\mathcal{T}}$ to be related to the two-point normalization of the displacement supermultiplet, as indeed is the case [29]. The relevant selection rules follows from the same argument presented in subsection 6.3 for (6.11), thus leading to
@@
and, from (7.3), we can extract the leading order data, thus obtaining
@@
and it is related by the mapping $z\rightarrow\chi$ to the other four-point function (7.3). As discussed in section 6, these correlation functions admit two distinct channels. For the chiral–antichiral expansion, the analysis is equivalent to the discussion presented above, whereas in the chiral–chiral channel one must consider the OPE [^15] See [22] for further details
@@
One can see that the quantum numbers of $L\bar{A}[1]_{1}^{(0)}$ are consistent with those of the displacement supermultiplet hence, in principle, it could appear in the exchange. However, this expansion must be related by an analogous relation of (6.9) to the one where the second and fourth insertions are swapped, i.e. braiding. Thus, consistency with (7.3) and (7.4) allows to exclude the superdisplacement in the possible exchanges. One can see this explicitly by considering the expansion up to NLO
@@
where, as discussed for (3.7), we collected all the contributions controlled by the three-point exchanges in $\text{h}^{(1)}(1-\chi)$. If now one considers the braiding relation for (7.5), one would obtain an extra term, related to $\text{h}^{(1)}(1-\chi)$, which is not present in (7.4). Thus leading to the conclusion that this extra term must vanish.
@@
by symmetry preserving argument. The other non-trivial three-point functions are given by
@@
Requiring invariance under the $\mathbb{Z}_{2}$ parity transformation $t\rightarrow-t$ imposes a selection rules on the parity of the third operator in the three-point function containing two fermions. This is essentially what has been discussed in (6.13), which applied to the present case leads to the expression
@@
This constrains the three-point function to be non-vanishing only for parity-odd contributions, as also observed in section 6. Specializing to the cases $\mathbb{T}_{\Delta}=\bar{\mathbbm{R}}$ and $\mathbb{T}_{\Delta}=\bar{\mathbbm{D}}$, which both have even parity, this immediately implies the vanishing of the corresponding three-point function coefficients, thus obtaining
@@
By recalling that these three-point coefficients also control the associated mixed multiparticle states, setting them to zero removes all such additional contributions, as discussed in 3.2.2. We thus find consistency with the results obtained from the braiding and we can conclude that the superdisplacement contributions are not exchanged.
It should be noted, however, that this discussion does not, in principle, exclude all supermultiplets with the quantum numbers of $L\bar{A}[1]_{1}^{(0)}$. Potential contributions associated with the composite operators of the tilt, encoded in $h^{(0)}(1-\chi)$ and $h^{(1)}(1-\chi)$, could in principle appear. Nevertheless, by arguments analogous to those discussed above, these contributions are absent, as they correspond to parity-even sectors, see section 6 and [21].
@@
Under similar considerations, this analysis could be extended to the study of all other four-point functions involving displacement, tilt and tlit multiplet. We do not pursue this here and simply note it as a possible direction for further study.
@@ Acknowledgments
I am grateful to Lorenzo Bianchi, Eduardo Casali and Andrea Mattiello for discussions and to Diego Trancanelli for discussions and comments on the draft. I would like to thank the organizers of the COST CA22113 meeting: "Recent Developments in Quantum Field Theory", where part of this work was presented and in particular Kiril Hristov also for hosting my visit at Sofia University. I thank COST Action CA22113 and BNSF, grant “Competition for Financial Support for Basic Research Projects – 2024”, for partial support.
@@
The 3-dimensional ${\mathcal{N}}=2$ superconformal algebra is $\mathfrak{osp}(2|4)$. Its bosonic subalgebra consists of the 3-dimensional conformal algebra $\mathfrak{so}(1,4)$ and of the R-symmetry algebra $\mathfrak{so}(2)_{R}$. The conformal generators are the rotations $M^{\mu\nu}$, the translations $P^{\mu}$, the special conformal transformations $K^{\mu}$ and the dilations $D$, with $\mu,\nu=0,1,2$ and algebra given by
@@
where, to avoid any confusion with spinorial indexes, we use an up left index convention for the particular combinations. In terms of $^{\pm}Q_{\alpha}$ and $^{\pm}S^{\alpha}$ the anti-commutation relations above now read
@@
The insertion of a 1/2 BPS Wilson line breaks the $\mathfrak{osp}(2|4)$ of the bulk theory down to $\mathfrak{su}(1,1|1)$. The $\mathfrak{su}(1,1)$ generators are those of the 1-dimensional conformal group, i.e. $\{D,P\equiv P_{0},K\equiv K_{0}\}$, satisfying
@@
As discussed, the R-symmetry ${\mathfrak{so}}(2)\simeq\mathfrak{u}(1)$ is fully preserved. As in [22] we choose $(\gamma^{\mu})_{\alpha}^{\;\;\beta}=(\sigma^{z},\sigma^{x},\sigma^{y})_{\alpha}^{\;\;\beta}$ as a basis and therefore $(\gamma^{\mu})_{\alpha\beta}=(\sigma^{1},-\sigma^{3},i\mathbb{1})$ and $(\gamma^{\mu})^{\alpha\beta}=(-\sigma^{1},\sigma^{3},i\mathbb{1})$. From the anticommutation relations one can identify the preserved fermionic charges, we are
@@
$[R,\bar{S}]$ $=-\bar{S},\qquad[M^{12},\bar{S}]=i\frac{1}{2}\bar{S}^{a},\qquad[J_{0},\bar{S}]=-\frac{1}{2}\bar{S}.$ (A.28)
@@ A.1 Blocks and superblocks
where the operators are taken as the sum, e.g. $D_{s}=D_{1}+D_{2}$ with $D_{i}$ acting on the $i$-th insertion, see [22, 44]. Given that only chargeless long multiplet are exchanged in this channel, the eigenvalue is $\mathfrak{c}_{\Delta,0}=\Delta^{2}$. By expanding in the Grassmann variables, one can identify the second-order differential equation

### v2

@@ 6.4 Universal sectors
We now turn to the conditions that characterize the presence of universal sectors. The analysis of the explicit OPEs shows that the displacement multiplet is never exchanged, so all the requirements discussed in Sec. 2 are satisfied. We therefore conclude that the superprimary and the displacement sectors remain decoupled up to NLO. The identification of the NLO four-point function then becomes straightforward, allowing us to bypass the standard bootstrap procedure of stating and fixing an ansatz. This case is very similar to that discussed in Sec. 4.1. The four-point function of the superprimary can be obtained directly from that of the first superdescendant of the ${\mathcal{N}}=4$ CSm Wilson line superprimary [52]. This leads to
@@
In this case, the overall coefficient is fixed explicitly by the analysis in Sec. 6.2 and the expansion parameter is again linked to the two-point normalization as in (4.8) by
@@ 6.5 Extracting the conformal data
6.5.1 Leading-order data
As discussed in Sec. 3 the exchanged operators at leading order are those of the multiparticle sates: $\mathbb{\Lambda}\bar{\mathbb{\Lambda}}]^{L=2}_{n}$ in the chiral-antichiral channels and $\mathbb{\Lambda}\mathbb{\Lambda}]^{L=2}_{n}$ in the chiral-chiral one, all having dimensions $\Delta=3+n$. We can thus extract the leading-order data by exploiting the orthogonal relation for the blocks and superblocks discussed in the Appendix A.1. We thus get
@@ 6.5.1 Leading-order data
6.5.2 Next-to-leading-order data
We now turn to the next-to-leading order of the expansion, beginning with the chiral-chiral channel. In this case, the anomalous dimensions are extracted from the logarithmic contributions of the NLO term in the correlation function (6.15). Explicitly, they are given by
@@ 6.5.2 Next-to-leading-order data
which explicitly reads
@@
Exploiting the relation under the map $z\rightarrow\chi$ discussed in Sec. 6.3 and the braiding relation (6.10), one can obtain the anomalous dimensions for the operators exchanged in the chiral-chiral channel. These are given by
@@
and, explicitly, the OPE coefficients read
@@ 7 The $1/3$ BPS Wilson line in ABJM
where the vertical lines stand for the action of the preserved supercharges corresponding to the considered theory and, for notation convenience, we also included the corresponding operators ($a=1,2,3$ and $r=1,2$). As in the order case, the four-point of the tilt and of the displacement (anti)chiral multiplet will be specified by the top components’ four-point functions of $\mathbb{R}$ ($\bar{\mathbb{R}}$) and $\mathbbm{F}$} ($\bar{\mathbbm{F}}$) respectively.
We now investigate the presence of the universal sector in both the displacement and tilt four-point functions. Focusing first on the displacement supermultiplet, one can verify exploiting the same arguments of Sec. 3 that the displacement sector does not mix with the tilt sector and, moreover, that the individual defect operators within the multiplet do not mix among themselves. It follows that the corresponding four-point function exhibits universal features. Furthermore, since the algebraic structure coincides with that of the $1/2$ BPS line in ${\mathfrak{osp}}(4|4)$, the analysis leads to the same expressions of [52].
@@
where $C_{\mathcal{T}}$ denotes the two-point normalization of the tilt supermultiplet, which is distinct from that of the displacement multiplet. Nevertheless, if the theory is controlled by a single perturbation parameter describing deviations from the GFF, one may still expect $C_{\mathcal{T}}$ to be related to the two-point normalization of the displacement supermultiplet, as indeed is the case [26]. The relevant selection rules follow from the same argument presented in Sec. 6.3 for (6.11), thus leading to
@@
and, from (7.3), we can extract the leading-order data, thus obtaining
@@
and it is related by the mapping $z\rightarrow\chi$ to the other four-point function (7.3). As discussed in Sec. 6, these correlation functions admit two distinct channels. For the chiral-antichiral expansion, the analysis is equivalent to the discussion presented above, whereas in the chiral-chiral channel one must consider the OPE [^16] See [52] for further details
@@
One can see that the quantum numbers of $L\bar{A}[1]_{1}^{(0)}$ are consistent with those of the displacement supermultiplet hence, in principle, it could appear in the exchange. However, this expansion must be related by an analogous relation of (6.9) to the one where the second and fourth insertions are swapped, i.e. braiding. Thus, consistency with (7.3) and (7.4) allows one to exclude the superdisplacement in the possible exchanges. One can see this explicitly by considering the expansion up to NLO
@@
where, as discussed for (3.7), we collected all the contributions controlled by the three-point exchanges in $\text{h}^{(1)}(1-\chi)$. If now one considers the braiding relation for (7.5), one would obtain an extra term, related to $\text{h}^{(1)}(1-\chi)$, which is not present in (7.4), and thus leading to the conclusion that this extra term must vanish.
@@
by symmetry preserving argument. The other nontrivial three-point functions are given by
@@
Requiring invariance under the $\mathbb{Z}_{2}$ parity transformation $t\rightarrow-t$ imposes a selection rule on the parity of the third operator in the three-point function containing two fermions. This is essentially what has been discussed in (6.13), which applied to the present case leads to the expression
@@
This constrains the three-point function to be nonvanishing only for parity-odd contributions, as also observed in Sec. 6. Specializing to the cases $\mathbb{T}_{\Delta}=\bar{\mathbbm{R}}$ and $\mathbb{T}_{\Delta}=\bar{\mathbbm{D}}$, which both have even parity, this immediately implies the vanishing of the corresponding three-point function coefficients, thus obtaining
@@
By recalling that these three-point coefficients also control the associated mixed multiparticle states, setting them to zero removes all such additional contributions, as discussed in Sec. 3.2.2. We thus find consistency with the results obtained from the braiding and we can conclude that the superdisplacement contributions are not exchanged.
It should be noted, however, that this discussion does not, in principle, exclude all supermultiplets with the quantum numbers of $L\bar{A}[1]_{1}^{(0)}$. Potential contributions associated with the composite operators of the tilt, encoded in $h^{(0)}(1-\chi)$ and $h^{(1)}(1-\chi)$, could in principle appear. Nevertheless, by arguments analogous to those discussed above, these contributions are absent, as they correspond to parity-even sectors; see Sec. 6 and [7].
@@
Under similar considerations, this analysis could be extended to the study of all other four-point functions involving displacement, tilt and tlit multiplets. We do not pursue this here and simply note it as a possible direction for further study.
@@ Acknowledgments
I am grateful to Lorenzo Bianchi, Eduardo Casali and Andrea Mattiello for discussions and to Diego Trancanelli for discussions and comments on the draft. I would like to thank the organizers of the COST CA22113 meeting: "Recent Developments in Quantum Field Theory", where part of this work was presented and in particular Kiril Hristov also for hosting my visit at Sofia University. I thank COST Action CA22113 and BNSF, grant “Competition for Financial Support for Basic Research Projects - 2024”, for partial support.
@@
The three-dimensional ${\mathcal{N}}=2$ superconformal algebra is $\mathfrak{osp}(2|4)$. Its bosonic subalgebra consists of the three-dimensional conformal algebra $\mathfrak{so}(1,4)$ and of the R-symmetry algebra $\mathfrak{so}(2)_{R}$. The conformal generators are the rotations $M^{\mu\nu}$, the translations $P^{\mu}$, the special conformal transformations $K^{\mu}$ and the dilations $D$, with $\mu,\nu=0,1,2$ and algebra given by
@@
where, to avoid any confusion with spinorial indexes, we use an up-left index convention for the particular combinations. In terms of $^{\pm}Q_{\alpha}$ and $^{\pm}S^{\alpha}$ the anti-commutation relations above now read
@@
The insertion of a 1/2 BPS Wilson line breaks the $\mathfrak{osp}(2|4)$ of the bulk theory down to $\mathfrak{su}(1,1|1)$. The $\mathfrak{su}(1,1)$ generators are those of the one-dimensional conformal group, i.e. $\{D,P\equiv P_{0},K\equiv K_{0}\}$, satisfying
@@
As discussed, the R-symmetry ${\mathfrak{so}}(2)\simeq\mathfrak{u}(1)$ is fully preserved. As in [52] we choose $(\gamma^{\mu})_{\alpha}^{\;\;\beta}=(\sigma^{z},\sigma^{x},\sigma^{y})_{\alpha}^{\;\;\beta}$ as a basis and therefore $(\gamma^{\mu})_{\alpha\beta}=(\sigma^{1},-\sigma^{3},i\mathbb{1})$ and $(\gamma^{\mu})^{\alpha\beta}=(-\sigma^{1},\sigma^{3},i\mathbb{1})$. From the anticommutation relations one can identify the preserved fermionic charges, which are
@@
$[R,\bar{S}]$ $=-\bar{S},\qquad[M^{12},\bar{S}]=i\frac{1}{2}\bar{S}^{a},\qquad[J_{0},\bar{S}]=-\frac{1}{2}\bar{S},$ (A.28)
@@ A.1 Blocks and superblocks
where the operators are taken as the sum, e.g. $D_{s}=D_{1}+D_{2}$ with $D_{i}$ acting on the $i$-th insertion; see [52, 36]. Given that only chargeless long multiplets are exchanged in this channel, the eigenvalue is $\mathfrak{c}_{\Delta,0}=\Delta^{2}$. By expanding in the Grassmann variables, one can identify the second-order differential equation